Jack can mow the yard in 4 hours. Ryan can mow it in 6 hours. how long will it take them to mow it together?
Answer by checkley77(12569)

(6*4)/(6+4)=24/10=2.4 HOURS TO MOW THE LAWN TOGETHER.